# Five parameters are all you need

### v1

@@ Abstract
$\Omega_{\mathrm{b}}$ $\in(0.04,0.06),$ $\quad\Omega_{\mathrm{m}}$ $\in(0.24,0.40),$ $\quad\zeta_{\mathrm{UV}}$ $\in(15,30).$ (6)
Their 1D and 2D projections in Figure 5 illustrate the sample uniformity in parameter space. Our 21cmFAST simulations have a 300 comoving Mpc box size and $768^{3}$ ($256^{3}$) cells for the matter (HI) field. We maintain most options in their default values and extract the neutral hydrogen fraction using lightcone.global_xH.
Figure 5: Sobol sampling of our $x_{\mathrm{HI}}$ profiles, in 6D parameter space of $\sigma_{8}$, $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{m}}$, and $\zeta_{\mathrm{UV}}$, Each point in the lower triangular panels corresponds to the 2D projection of a 21cmFAST run, while the diagonal panels show the 1D cumulative histograms for each parameter. These 1D and 2D projections demonstrate the uniformity of our sampling of the parameter space within the prior range of Simulations.
The ionization efficiency $\zeta_{\mathrm{UV}}$ governs the ability of ultraviolet photons to escape their parent galaxies and ionize the IGM. An increase in $\zeta_{\mathrm{UV}}$ leads to an earlier completion of HI reionization. Due to significant uncertainties surrounding $\zeta_{\mathrm{UV}}$, we opted to use a constant value in each simulation. Note that this choice was made for simplicity, and a more realistic assumption could be that $\zeta_{\mathrm{UV}}$ is a function of halo mass 53 or redshift. However, our findings suggest that $\zeta_{\mathrm{UV}}$ plays a minor role compared to the cosmological parameters. In fact, the best performing symbolic regression expressions (see Symbolic regression for the selection of best models) do not use $\zeta_{\mathrm{UV}}$.
@@
The early intergalactic medium is primarily composed by neutral hydrogen and helium. Neutral helium (HeI) loses is first electron at the same time as neutral hydrogen (HI) gets ionized 54. However, there is a second reionization that occurs around $z\sim 3$ where Helium (HeII) loses its remaining electron.
CMB photons will scatter off any free electrons, therefore both helium reionizations contribute to the Thomson optical depth to reionization $\tau_{\mathrm{reio}}$, although HeII ionization contribute relatively little in comparison to HI and HeI ionizations 55.
@@
Given the relatively small impact of the HeII reionization on $\tau_{\mathrm{reio}}$, the current uncertainties regarding its timeline, and the difficulty involved with its accurate modeling 56, 57, we opt to follow the conventional approach and include the second Helium reionization using the $\tanh$ model
@@
Figure 6: Universal shape of $x_{\mathrm{HI}}$. (Top) 128 simulated $x_{\mathrm{HI}}$ timelines (thin light gray lines) exhibit universality, after power-law transformations $\tilde{a}=(a/\alpha)^{\beta}$. The blue curves in all panels show our fitted analytic shape model, a composition of the Gompertz curve with a 5th-degree polynomial in Eqs. 2 and 3. (Middle) Time derivative of $x_{\mathrm{HI}}$, can be interpreted as a PDF if we view $x_{\mathrm{HI}}$ itself as the CDF. We first discovered the universality by translating and rescaling each $x_{\mathrm{HI}}$ using the mean and variance of its PDF, though now switch to the better approach that jointly fits the global shape and individual power-law parameters. We use the latter as target of symbolic regression. (Bottom) Timelines transformed by the inverse of Gompertz function, modeled in blue curve with a 5th-degree polynomial in Eq. 3.
@@
To address this, we construct flexible models for the universal shape, and fit it jointly with individual transformation parameters of each $x_{\mathrm{HI}}$ timeline. We compose the Gompertz function $\mathrm{gomp}$, defined in Eq. 2, with a low-degree polynomial $P_{m}$, where $m=1,3,5,7$ progressively. We fit the composed shape to minimize the mean squared error (MSE) in 128 $x_{\mathrm{HI}}$’s and at 92 time points in each, and find the objective value improve with $m$ but only marginally from $P_{5}$ to $P_{7}$. Therefore, our final shape model is a composition of $\mathrm{gomp}$ and $P_{5}$, (see the lower panel of Figure 6), and has 6 parameters to fit.
As for the transformation parameters, as in the PDF approach, we use an affine transformation $\ln\tilde{a}=\beta(\ln a-\ln\alpha)$, or equivalently a power law in Eq. 4. Because each $x_{\mathrm{HI}}$ has its own parameters of $\alpha$ and $\beta$, we have in total $262=6+2\times 128$ parameters to determine in the joint fit. Figure 6 shows all 128 $x_{\mathrm{HI}}$ timelines and their universality after transformations. We can then use the fitted $\alpha$ and $\beta$ as the target for symbolic regression, to model their cosmological dependences on the other independent parameters.
@@
Figure 7: Pareto front for symbolic regression (blue solid steps) illustrates the trade-off between regression accuracy and expression complexity. We further add the lower convex hull to aid model selection: each segment of the orange dotted line represents a power-law trade-off in this log-log plot, and every model that touches the convex hull is more economic – in the sense of accuracy gain at cost of complexity – than the nearby models above the segments. As an example, here we show the results of $\ln\alpha(\bm{\theta})$, where the complexity 11 point gives Eq. 5.
SR learns a model of data in the form of analytic expressions. Unlike traditional fittings that are restricted by their specific parameterization, SR searches in the vast function space of all expressions composed of specified operations, input variables, and free constants. It is NP-hard 58, 59 and typically need genetic or deep learning-based algorithms. In this work, we use the PySR package 60, 21 which performs SR optimizations using a multi-population genetic algorithm. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, it’d be great if you can refine/rewrite the SR and PySR intro
With PySR, we search for symbolic expressions that take the 6 parameters in Figure 5 as inputs and output $\ln\alpha$ or $\beta$, to minimize the MSE loss. Here, the search space is the set of expressions composed of 5 binary operators ($+$, --, $\cdot$, ///, and power function) and 2 unary operators ($\exp$ and $\ln$). Each expression naturally takes the form of a binary tree, and the total number of nodes is a measure of its complexity. We use 512 PySR populations each having 33 expressions, and optimize for 10000 iterations each with 10000 cycles. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, could you translate this into some brief description in the language of evolve-simplify-optimize loop, mutation, crossover, and tournament?
@@
For the pivot $\ln\alpha$, we find complexity 11 is enough for the MSE loss to reach $10^{-4}$ (sub-percent level error on average), so there’s no need to use expressions more complex that that. However, for the tilt $\beta$, complexity 1, i.e. a constant, fit with a loss $\approx 0.1$ ($4\%$ error on average), while higher complexities can help but very slowly. Therefore, based on the economic heuristic, we choose expressions of complexity 11 and 1, for $\ln\alpha$ and $\beta$, respectively.
This preference for constant tilt could be a consequence of the astrophysical assumptions of reionization in 21cmFAST. A more complex simulation, such as one involving radiative transfer or different X-ray preheating 61, might reveal a different tilt.
@@
Gompertz reionization Tanh reionization Parameter gomp 0 gomp 1 gomp 2 $\tanh$ 0 $\tanh$ 1 Planck PR31 $10^{9}A_{\mathrm{s}}$ 2.101(31) 2.088(12) 2.089(13) 2.099(30) 2.098(30) 2.100(30) $n_{\mathrm{s}}$ 0.9638(41) 0.9634(40) 0.9634(40) 0.9641(42) 0.9641(41) 0.9649(42) $\Omega_{\mathrm{c}}h^{2}$ 0.1201(12) 0.1203(11) 0.1202(11) 0.1201(12) 0.1201(12) 0.1200(12) $\Omega_{\mathrm{b}}h^{2}$ 0.02235(15) 0.02233(14) 0.02233(14) 0.02235(15) 0.02235(15) 0.02237(15) $\Omega_{\mathrm{m}}$ 0.3157(76) 0.3171(68) 0.3168(68) 0.3157(76) 0.3159(75) 0.3153(73) $\Omega_{\mathrm{m}}h^{2}$ 0.1430(11) 0.1433(10) 0.1432(10) 0.1430(11) 0.1431(11) 0.1430(11) $\Omega_{\Lambda}$ 0.6842(76) 0.6828(68) 0.6831(68) 0.6842(76) 0.6840(75) 0.6847(73) Age 13.800(23) 13.803(20) 13.803(22) 13.800(23) 13.800(23) 13.797(23) $H_{0}$ 67.32(55) 67.22(49) 67.24(50) 67.32(55) 67.31(54) 67.36(54) $100\theta_{\mathrm{X}}$ 1.04184(29) 1.04183(29) 1.04183(29) 1.04815(29) 1.04184(29) 1.04092(31) $\sigma_{8}$ 0.8109(59) 0.8092(45) 0.8092(46) 0.8108(60) 0.8105(59) 0.8111(60) $S_{8}$ 0.832(13) 0.832(13) 0.832(13) 0.832(13) 0.832(13) 0.832(13) $\tau_{\mathrm{reio}}$ 0.0546(76) 0.05115(61) 0.05140(63) 0.0543(75) 0.0538(74) 0.0544(73) $z_{\mathrm{re}}$ 6.81(68) 7.40 7.41 7.67(75) 7.62(74) 7.67(73)
Figure 8: Analysis of CMB data treating $\tau_{\mathrm{reio}}$ as a derived parameter using Eqs. 2, 3, 4 and 5 vs. sampling it with the conventional $\tanh$ model. Here gomp 1 combines our Gompertz universal shape with the rescaling pivot and tilt obtained via symbolic regression. Note that for gomp 1 we do not sample over $z_{\mathrm{re}}$ since Eq. 2 does not depend on it. The green (blue) contours correspond to the constraints obtained with our Gompertz universal shape ($\tanh$ 1 model).
Figure 9: Validation of our universal shape for $x_{\mathrm{HI}}$ in Eqs. 2 and 3 vs. the conventional $\tanh$ model using CMB data with sampling over $\tau_{\mathrm{reio}}$. Note that we have not eliminated $\tau_{\mathrm{reio}}$ from the analysis yet. Here, we use the standard choice of sampling over optical depth and obtain the corresponding reionization history via bisection. The green (blue) contours correspond to the constraints obtained with our gomp 0 ($\tanh$ 0) model.
@@
To investigate the impact of overfitting, we train an alternative mapping of reionization with cosmology using only half of the 21cmFAST $x_{\mathrm{HI}}$ profiles, which we refer as gomp 2. For the tilt, once again we find a constant $\beta=8.331$, remarkably close to the value obtained using the full simulated data (8.290). And for the pivot, we obtain
@@
$\ln\alpha(\bm{\theta})=(1.039+\sigma_{8})(\Omega_{\mathrm{b}}-\Omega_{\mathrm{m}}h-n_{\mathrm{s}}),$ (10)
with a training loss of $1.2\times 10^{-4}$. Now we can test this on the other half of our $x_{\mathrm{HI}}$ sample and get a validation loss of $1.4\times 10^{-4}$, only slightly larger than the training loss implying we are safe from overfitting.
Unsurprisingly, Eq. 10 recovers the cosmological trends present in Eq. 5. Specifically, increasing $n_{\mathrm{s}}$, $\Omega_{\mathrm{m}}$, and $\sigma_{8}$ leads to an earlier reionization scenario, while a higher value of $\Omega_{\mathrm{b}}$ slightly delays reionization. Furthermore, the impact of the Hubble constant is once again linked to the matter density, albeit not in the same manner as in Eq. 5. This is likely evidence that the role of $h$ is limited, and PySR considers its influence in conjunction with $\Omega_{\mathrm{m}}$ without increasing the complexity of the analytic expression.
@@
Replacing Eq. 5 with Eq. 10, we rerun the analysis following the same approach, and summarize the results in Table 1. Comparing these findings to our previous results, we observe strikingly similar values for all cosmological parameters, albeit with smaller deviations from the Planck results. The consistency between our results using the full $x_{\mathrm{HI}}$ sample and those using only half of it indicates that our symbolic expression mappings perform robustly and are not biasing the parameter inferences.
@@
Table 1 summarizes the results obtained by performing MCMC Bayesian inference with Cobaya for the models considered throughout this work. We also include the Planck results 1 for reference. In total, we run three Gompertz reionization and two $\tanh$ reionization models, where $\tanh$ 0, Planck PR3, and gomp 0 sample over the typical 6 cosmological parameters (see shaded parameters in Table 1). In contrast, gomp 1, and gomp 2 sample over $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},\Omega_{\mathrm{b}}h^{2},\Omega_{\mathrm{c}}h^{2},h\}$. Similarly, $\tanh$ 1 samples over the same parameters as gomp 1 but with the inclusion of $z_{\mathrm{re}}$, and agrees with $\tanh$ 0 as the expected result from this sanity check.
@@
There is an apparent discrepancy between the different models in $\theta_{\mathrm{x}}$, a proxy for the angular scale of the acoustic oscillations ($\theta_{*}$). The difference is due to our use of $100\theta_{\mathrm{s}}$, which corresponds to the peak scale parameter defined exactly as the ratio of the sound horizon divided by the angular diameter at decoupling, with decoupling time given by the maximum of the visibility function, i.e. the standard choice for CLASS. In contrast, the Planck collaboration reports $100\theta_{\mathrm{MC}}$, which is given by Eq. (6) in 62, and corresponds to the standard choice in CosmoMC26.
@@
Although Table 1 does not display an error in the midpoint of reionization in the gomp 1 and gomp 2 models, it is important to note that an error does exist. Unfortunately, it cannot be automatically computed with Cobaya because gomp 1 and gomp 2 do not sample over reionization parameters. Therefore, $z_{\mathrm{re}}$ has no meaning in our Gompertz CLASS [^3] This oversight can be corrected; however, it would require re-running the chains. for models that take advantage of Eq. 1.
@@
Remarkably, the inferred value of $S_{8}$ remains consistent across all reionization models, even when $\sigma_{8}$ and $\Omega_{\mathrm{m}}$ vary between models. This trend suggests that different reionization models neither alleviate nor exacerbate the $S_{8}$ tension. While this observation holds for the Planck PR3 data, we anticipate that incorporating low-redshift Baryon Acoustic Oscillation (BAO) data may affect this conclusion. Future work will explore the implications of Gompertz reionization for the joint analysis of CMB and low-redshift data.

### v2

@@ Abstract
$\Omega_{\mathrm{b}}$ $\in(0.04,0.06),$ $\quad\Omega_{\mathrm{m}}$ $\in(0.24,0.40),$ $\quad\zeta_{\mathrm{UV}}$ $\in(20,35).$ (6)
[…]
Our 21cmFAST simulations have a 300 comoving Mpc box size and $768^{3}$ ($256^{3}$) cells for the matter (HI) field. We maintain most options in their default values and extract the neutral hydrogen fraction using lightcone.global_xH.
Figure 5: Sobol design for our $x_{\mathrm{HI}}$ edge and core samples, in 6D parameter space of $\sigma_{8}$, $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{m}}$, and $\zeta_{\mathrm{UV}}$. Each of the 128 points in the lower triangular panels corresponds to the 2D projection of one core and one edge 21cmFAST runs, while the diagonal panels show the 1D cumulative histograms for each parameter. These 1D and 2D projections demonstrate the uniformity of our sampling of the parameter space within the edge prior range of Simulations, as well as in the core area surrounding the Planck PR3 constraint.
The ionization efficiency $\zeta_{\mathrm{UV}}$ governs the ability of ultraviolet photons to escape their parent galaxies and ionize the IGM. An increase in $\zeta_{\mathrm{UV}}$ leads to an earlier completion of HI reionization. Due to significant uncertainties surrounding $\zeta_{\mathrm{UV}}$, we opted to use a constant value in each simulation. Note that this choice was made for simplicity, and a more realistic assumption could be that $\zeta_{\mathrm{UV}}$ is a function of halo mass 46 or redshift, which we will investigate in future work.
@@
The early intergalactic medium is primarily composed by neutral hydrogen and helium. Neutral helium (HeI) loses is first electron at the same time as neutral hydrogen (HI) gets ionized 47. However, there is a second reionization that occurs around $z\sim 3$ where Helium (HeII) loses its remaining electron.
CMB photons will scatter off any free electrons, therefore both helium reionizations contribute to the Thomson optical depth to reionization $\tau_{\mathrm{reio}}$, although HeII ionization contribute relatively little in comparison to HI and HeI ionizations 48.
@@
Given the relatively small impact of the HeII reionization on $\tau_{\mathrm{reio}}$, the current uncertainties regarding its timeline, and the difficulty involved with its accurate modeling 49, 50, we opt to follow the conventional approach and include the second Helium reionization using the $\tanh$ model
@@
Figure 6: Universal shape of $x_{\mathrm{HI}}$. (Top) 256 simulated $x_{\mathrm{HI}}$ timelines (thin light gray lines) exhibit universality, after power-law transformations $\tilde{a}=(a/\alpha)^{\beta}$. The blue curves in all panels show our fitted analytic shape model, a composition of the Gompertz curve with a 5th-degree polynomial in Eqs. 1 and 2. (Middle) Time derivative of $x_{\mathrm{HI}}$, can be interpreted as a PDF if we view $x_{\mathrm{HI}}$ itself as the CDF. We first discovered the universality by translating and rescaling each $x_{\mathrm{HI}}$ using the mean and variance of its PDF, though now switch to the better approach that jointly fits the global shape and individual power-law parameters. We use the latter as target of symbolic regression. (Bottom) Timelines transformed by the inverse of Gompertz function, modeled in blue curve with a 5th-degree polynomial in Eq. 2.
@@
To address this, we construct flexible models for the universal shape, and fit it jointly with individual transformation parameters of each $x_{\mathrm{HI}}$ timeline. We compose the Gompertz function $\mathrm{gomp}$, defined in Eq. 1, with a low-degree polynomial $P_{m}$, where $m=1,3,5,7$ progressively. We fit the composed shape to minimize the mean squared error (MSE) in 256 $x_{\mathrm{HI}}$’s and at 127 time points in each, and find the objective value improve with $m$ but only marginally from $P_{5}$ to $P_{7}$. Therefore, our final shape model is a composition of $\mathrm{gomp}$ and $P_{5}$, (see the lower panel of Figure 6), and has 6 parameters to fit.
As for the transformation parameters, as in the PDF approach, we use an affine transformation $\ln\tilde{a}=\beta(\ln a-\ln\alpha)$, or equivalently a power law in Eq. 3. Because each $x_{\mathrm{HI}}$ has its own parameters of $\alpha$ and $\beta$, we have in total $518=6+2\times 256$ parameters to determine in the joint fit. Figure 6 shows all 256 $x_{\mathrm{HI}}$ timelines and their universality after transformations. We can then use the fitted $\alpha$ and $\beta$ as the target for symbolic regression, to model their dependences on the independent cosmological and astrophysical parameters.
@@
Figure 7: Pareto front for symbolic regression (blue solid steps) illustrates the trade-off between regression accuracy and expression complexity. We further add the lower convex hull to aid model selection: each segment of the orange dotted line represents a power-law trade-off in this log-log plot, and every model that touches the convex hull is more economic – in the sense of accuracy gain at cost of complexity – than the nearby models above the segments. As an example, here we show the results of $\ln\alpha(\bm{\theta})$, where the complexity 22 point gives Eq. 4.
SR learns a model of data in the form of analytic expressions. Unlike traditional fittings that are restricted by their specific parameterization, SR searches in the vast function space of all expressions composed of specified operations, input variables, and free constants. It is NP-hard 51, 52 and typically need genetic or deep learning-based algorithms. In this work, we use the PySR package 53, 25 which performs SR optimizations using a multi-population genetic algorithm. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, it’d be great if you can refine/rewrite the SR and PySR intro
With PySR, we search for symbolic expressions that take the 6 parameters in Figure 5 as inputs and output $\ln\alpha$ or $\beta$, to minimize the MSE loss. Here, the search space is the set of expressions composed of 5 binary operators ($+$, --, $\cdot$, ///, and power function) and 2 unary operators ($\exp$ and $\ln$). Each expression naturally takes the form of a binary tree, and the total number of nodes is a measure of its complexity. We use 512 PySR populations each having 33 expressions, and optimize for 20000 iterations each with 10000 cycles. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, could you translate this into some brief description in the language of evolve-simplify-optimize loop, mutation, crossover, and tournament?
@@
Our current selection for tilt and pivot could be a consequence of the astrophysical assumptions of reionization in 21cmFAST. A more complex simulation, such as one involving radiative transfer or different X-ray preheating 54, might reveal a different mapping between physical parameters and $x_{\mathrm{HI}}$ profiles.
@@
To investigate the impact of overfitting, we train an alternative mapping of reionization with cosmology using only half of the 21cmFAST $x_{\mathrm{HI}}$ profiles, which we refer as SRHalf. For the pivot and tilt, we obtain
@@
$\ln\alpha(\bm{\theta})$ $=\Bigl{(}\frac{\Omega_{\mathrm{b}}}{\Omega_{\mathrm{m}}}\Bigr{)}^{\Omega_{\mathrm{m}}}-\ln^{0.5271}\Bigl{(}h\bigl{(}\zeta_{\mathrm{UV}}+\Omega_{\mathrm{b}}^{-0.4982}\bigr{)}^{\sigma_{8}}\Bigr{)}-n_{\mathrm{s}}^{1.834},$ (10)
$\beta(\bm{\theta})$ $=\biggl{(}\frac{\zeta_{\mathrm{UV}}-\Omega_{\mathrm{m}}^{-1.583}}{\Omega_{\mathrm{b}}h}\biggr{)}^{0.3163},$ (11)
with a training loss (complexity) of $1.1\times 10^{-5}$ (22) and $3.7\times 10^{-3}$ (11), respectively. Overall, we obtain a relative root mean squared error of 0.9% when reproducing the simulated optical depths from the $x_{\mathrm{HI}}$ (edge + core) with SRHalf. Now we can test this on the other half of our $x_{\mathrm{HI}}$ sample and get a validation loss of $2.2\times 10^{-5}$ and $6.5\times 10^{-3}$ respectively, only slightly larger than the training loss on absolute scales, which implies that we are safe from overfitting, supported by consistency of their MCMC results.
Unsurprisingly, Eqs. 10 and 11 recover the cosmological trends discovered with SRFull. Specifically, increasing $\zeta_{\mathrm{UV}}$, $n_{\mathrm{s}}$, $\Omega_{\mathrm{m}}$, and $\sigma_{8}$ primarily leads to earlier reionization, while larger $\Omega_{\mathrm{b}}$ delays it. The ratio of matter densities appears in the pivot expression again. Notably, the role of $h$ generally shows the opposite trend in $\alpha$. Furthermore, Eq. 11 excludes $n_{\mathrm{s}}$ and $\sigma_{8}$ because of our choice of complexity for SRHalf, which, despite having a larger loss, still achieves mean squared errors comparable to SRFull when reproducing the simulated data.
@@
Replacing Eqs. 4 and 5 with Eqs. 10 and 11, we rerun the analysis following the same approach, and summarize the results using CMB jointly with astrophysical data in Table 1. Comparing these findings to our previous results, we observe strikingly similar values for all cosmological parameters. The consistency between our results using the full $x_{\mathrm{HI}}$ sample and those using only half of it indicates that our symbolic expression mappings perform robustly and are not biasing the parameter inferences.
Astrophysical data
[…]
In this work, we supplement Planck CMB data with quasar damping wing (DW) constraints on $x_{\mathrm{HI}}$ (green pentagons in Figure 4), reflecting our confidence in their robustness. Future studies could include luminosity function (LF) constraints as more data becomes available, particularly since we find including them could already halve the error on $\tau_{\mathrm{reio}}$, and because more data will enhance our ability to constrain and marginalize over reionization astrophysics. We opt not to include the LF constraints on $x_{\mathrm{HI}}$ in this work as they are in slight tension with the DW data at the current stage (Figure 4).
[…]
CMB CMB + DW Parameter Planck PR31 $\tanh$ gomp gomp + SRFull gomp + SRHalf $10^{9}A_{\mathrm{s}}$ 2.100(30) 2.099(30) 2.101(31) 2.094(13) 2.094(13) $n_{\mathrm{s}}$ 0.9649(42) 0.9641(42) 0.9640(42) 0.9636(39) 0.9637(39) $\Omega_{\mathrm{c}}h^{2}$ 0.1200(12) 0.1201(12) 0.1200(12) 0.1202(11) 0.1202(11) $\Omega_{\mathrm{b}}h^{2}$ 0.02237(15) 0.02235(15) 0.02235(15) 0.02234(14) 0.02234(14) $\Omega_{\mathrm{m}}$ 0.3153(73) 0.3157(76) 0.3155(75) 0.3166(69) 0.3163(69) $\Omega_{\mathrm{m}}h^{2}$ 0.1430(11) 0.1430(11) 0.1430(11) 0.1432(10) 0.1431(10) ) $\Omega_{\Lambda}$ 0.6847(73) 0.6842(76) 0.6845(75) 0.6833(69) 0.6836(69) Age 13.797(23) 13.800(23) 13.799(23) 13.802(22) 13.801(22) $H_{0}$ 67.36(54) 67.32(55) 67.34(55) 67.26(50) 67.28(50) $100\theta_{\mathrm{X}}$ 1.04092(31) 1.04185(29) 1.04185(29) 1.04184(29) 1.04184(29) $\sigma_{8}$ 0.8111(60) 0.8108(60) 0.8109(60) 0.8100(44) 0.8099(44) $S_{8}$ 0.832(13) 0.832(13) 0.832(13) 0.832(13) 0.832(13) $\tau_{\mathrm{reio}}$ 0.0544(73) 0.0543(75) 0.0547(76) 0.0526(15) $\mathbf{0.0527^{+(14)}_{-(16)}}$ $z_{\mathrm{re}}$ 7.67(73) 7.67(75) 6.76(67) 6.98 6.99 $\zeta_{\mathrm{UV}}$ $\mathbf{26.9^{+2.1}_{-2.5}}$ $\mathbf{27.0^{+2.1}_{-2.5}}$
@@
Our inference combines CMB with astrophysical data. Table 1 summarizes the results obtained by performing MCMC Bayesian inference with Cobaya for the models considered throughout this work. We also include the Planck results 1 for reference. In total, we run one Gompertz and one $\tanh$ reionization models with CMB data and two extra Gompertz models using SR and astrophysical data jointly with CMB to infer cosmology while constraining and marginalizing over $\zeta_{\mathrm{UV}}$.
When considering only CMB data, we sample the typical 6 cosmological parameters (see shaded parameters in Table 1), sampling $\tau_{\mathrm{reio}}$ and using bisection to obtain the reionization timeline. For $\tanh$, the bisection varies $z_{\mathrm{re}}$, while for Gompertz we fix $\beta$ to the mean value (7.66, the complexity 1 expression from PySR) and vary $\alpha$. Therefore, we do not use the SR mapping from physical parameters to reionization – Eqs. 4 and 5 or alternatively Eqs. 10 and 11.
In contrast, when astrophysical data is included, $\tanh$ samples over $z_{\mathrm{re}}$ instead of $\tau_{\mathrm{reio}}$ and an additional likelihood component is calculated in Cobaya. For Gompertz, we do not sample any reionization parameter. Instead, we use SRFull or SRHalf to map physical parameters to reionization timeline and include the astrophysical likelihood component. Perhaps unsurprisingly, $\tanh$ reionization does a poor job of fitting the astrophysical data, with a $\chi^{2}$ larger than Gompertz’s by an order of magnitude, hence we do not show the $\tanh$ + CMB + DW results in Table 1.
@@
There is an apparent discrepancy between the different models in $\theta_{\mathrm{x}}$, a proxy for the angular scale of the acoustic oscillations ($\theta_{*}$). The difference is due to our use of $100\theta_{\mathrm{s}}$, which corresponds to the peak scale parameter defined exactly as the ratio of the sound horizon divided by the angular diameter at decoupling, with decoupling time given by the maximum of the visibility function, i.e. the standard choice for CLASS. In contrast, the Planck collaboration reports $100\theta_{\mathrm{MC}}$, which is given by Eq. (6) in 65, and corresponds to the standard choice in CosmoMC31. Unsurprisingly, different reionization scenarios give similar values of $100\theta_{\mathrm{s}}$.
@@
Although Table 1 does not display an error in the midpoint of reionization for gomp + SRFull + CMB + DW nor gomp + SRHalf + CMB + DW models, it is important to note that an error does exist. Unfortunately, it cannot be automatically computed with Cobaya because SRFull and SRHalf models do not sample over reionization parameters. Therefore, $z_{\mathrm{re}}$ has no meaning in our Gompertz CLASS [^2] This oversight can be corrected; however, it would require re-running the chains. for models that take advantage of the mapping between physical parameters and reionization timeline.
@@
Remarkably, the inferred value of $S_{8}$ remains consistent across all reionization models, even when $\sigma_{8}$ and $\Omega_{\mathrm{m}}$ show small variations between models. This trend suggests that different reionization models neither alleviate nor exacerbate the $S_{8}$ tension. While this observation holds for the Planck PR3 data and for CMB + DW, we anticipate that incorporating low-redshift Baryon Acoustic Oscillation (BAO) data may affect this conclusion. Future work will explore the implications of Gompertz reionization for the joint analysis of CMB and low-redshift data.
Figure 8: Analysis of CMB data and astrophysical data treating $\tau_{\mathrm{reio}}$ as a derived parameter using Eqs. 1, 2, 3, 4 and 5 vs. sampling it with the conventional $\tanh$ model. Here gomp + SRFull combines our Gompertz universal shape with the rescaling pivot and tilt obtained via symbolic regression. Note that for gomp + SRFull we do not sample over $z_{\mathrm{re}}$, since Eq. 1 does not depend on it but we use the astrophysical data to constrain and marginalize over reionization astrophysics. The green (blue) contours correspond to the constraints obtained with our Gompertz universal shape ($\tanh$ model).
Figure 9: Validation of our universal shape for $x_{\mathrm{HI}}$ in Eqs. 1 and 2 vs. the conventional $\tanh$ model using CMB data with sampling over $\tau_{\mathrm{reio}}$. Note that we have not use our mapping from physical parameters to $\tau_{\mathrm{reio}}$. Here, we use the standard choice of sampling over optical depth and obtain the corresponding reionization history via bisection. The green (blue) contours correspond to the constraints obtained with our gomp ($\tanh$) model.